OpenStudy (anonymous):

What is -0.75 as a fraction in simplest form?

OpenStudy (kainui):

Remember that for every amount it is past the decimal, you are adding a 0. For instance: .5 is in the tenths place and we can rewrite it as 5/10. Similarly, .05 is in the hundredths place, so we write it as 5/100. However these aren't simplified, you need to divide out common factors, so if we look at 5/10 we can easily see that: \[\LARGE \frac{5}{10}=\frac{5}{5*2}=\frac{5}{5} *\frac{1}{2}=\frac{1}{2}\] since anything divided by itself is just 1, the 5's cancel each other out.